OpenStudy (anonymous):

Complete the given sequence: 0, 5, 8, 17, ___

OpenStudy (anonymous):

ahhhhh nvm sorry thought it was 3^2, 3^3, 3^4 :P sorry nvm then i saw the 0, 5 :P

OpenStudy (anonymous):

wtf??

OpenStudy (anonymous):

n^2 + 1

OpenStudy (anonymous):

sathipraskash... I am sorry I gave you the answer... I don't appreciate the wtf

OpenStudy (anonymous):

no nt for u..

OpenStudy (anonymous):

would u plz explain blex....

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Well.. it isn't what I thought. I think it is alternating terms 1^2 - 1 = 0 2^2 + 1 = 5 3^2 - 1 = 8 4^2 + 1 = 17 5^2 - 1 = 24 6^2 + 1 = 37 7^2 - 1 = 48 8^2 + 1 = 65 etc.